place one set of parenthesis on the left side of the equation to make it true
32+4-12/2*8-5+12/3+7=13
thanks for any help this is driving me nuts

Answer by sudhanshu_kmr(1152)


32+4-12/2*(8-5+12)/3+7
= 32 + 4-12/2 * 15/3 +7
= 32 + 4 - 6 * 5 +7
= 32 + 4 -30 +7
= 13